(6x^5y^2z^-4)/(4x^-2y^4z^-1)
Answer by user_dude2008(1862)

(6x^5y^2z^(-4))/(4x^(-2)y^4z^(-1)) = (3x^(5--2)y^(2-4)z^(-4--1))/2 = (3x^(5+2)y^(2-4)z^(-4+1))/2 = (3x^(7)y^(-2)z^(-3))/2 = (3x^7)/(2y^2z^3)


Answer: (3x^7)/(2y^2z^3)